OpenStudy (anonymous):

find the length: f(x) = x^2-1/8 ln x, 1

OpenStudy (anonymous):

take the derivative square it then add one those are the first three steps, mostly algebra usually the problem is cooked up to be easy

OpenStudy (anonymous):

is it \[f(x)=\frac{x^2-1}{8}\]?

OpenStudy (anonymous):

no its f(x) = x^2 - (1/8) ln x

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

then the derivative is \(2x-\frac{1}{8x}\) i guess right?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

then square it

OpenStudy (anonymous):

4x^2+1/64x^2-1/2

OpenStudy (anonymous):

that is the annoying algebra part ok then add one

OpenStudy (anonymous):

it would be +1/2 instead of -1/2

OpenStudy (anonymous):

then write as a single fraction

OpenStudy (anonymous):

uhhhhh

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

gotta add no way around it

OpenStudy (anonymous):

4x^2 + 1/64x^2 +1/2 ?!

OpenStudy (anonymous):

the denominator will be \(64x^2\) the only denominator you see the numerator will be \[4x^2\times 62x^2+32x^2+1\]

OpenStudy (anonymous):

uhhhhhhhhhhh that blew my mind lol

OpenStudy (anonymous):

nay it is how you add is all what should blow your mind is the next step

OpenStudy (anonymous):

uhh whats the next step O_o

OpenStudy (anonymous):

the miracle that the numerator is actually a perfect square you have to admire the work that when it to this problem the numerator is .... \[(16x^2+1)^2\]

OpenStudy (anonymous):

btw the algebra to add is not that bad at all, i am sure you know how to add to recognize it as a perfect square in the ingeneos part

OpenStudy (anonymous):

yeah i get it till here now

OpenStudy (anonymous):

however you spell it and it has to be a perfect square because the next step it to take the square root!!

OpenStudy (anonymous):

then multiply by 2 pi x

OpenStudy (anonymous):

your final job is to take the square root, and integrate the square root of a square is easy enough you get \[\int_a^b\frac{16x^2+1}{8x}dx\]

OpenStudy (anonymous):

nah it is the arc length right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

oh yeah i thought it was surface area sorry

OpenStudy (anonymous):

thank god it is \[\int_a^b\sqrt{1+(f')^2}dx\]

OpenStudy (anonymous):

imagine all the work that went in to making \[1+(f')^2\] a perfect square i tried it once and it took the better part of an hour

OpenStudy (anonymous):

oh damnnnnn. i just hope he don't put any on the exam

OpenStudy (anonymous):

tomorrow

OpenStudy (anonymous):

maybe a real easy one or just "set up but don't evaluate" anyway i hope the steps are clear the integral is now routine good luck on your exam

OpenStudy (anonymous):

so i take the integral of that the one you sent

OpenStudy (anonymous):

thank you so much for the help.

OpenStudy (anonymous):

yeah that is it if forgot the limits

OpenStudy (anonymous):

break it in to two parts the first is easy, the second is the log, that is why one limit had an \(e\) in it i guess

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

take 8x out ?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

oh sorry i was thinking something else lol

OpenStudy (anonymous):

\[\frac{16x^2+1}{8x}=2x+\frac{1}{8x}\]

OpenStudy (anonymous):

take 8x as u

OpenStudy (anonymous):

integral is therefore \[x^2+\frac{1}{8}\ln(x)\]

OpenStudy (anonymous):

you are thinking way way too hard

OpenStudy (anonymous):

oh its already in dx form

OpenStudy (anonymous):

so the answer is the one you gave?

OpenStudy (anonymous):

that is the integral, yes convince yourself it was a real real easy one then evaluate etc good luck

OpenStudy (anonymous):

now just plug the values?

OpenStudy (anonymous):

1 and e and f(B)-f(A)

OpenStudy (anonymous):

wait where did the 8x came from in the denominator

OpenStudy (tkhunny):

\(f(x) = x^{2}-(1/8) ln x,\;1<x<e\) \(f'(x) = 2x - 1/(8x) = \dfrac{16x^2 - 1}{8x},\;1<x<e\)